# Adversarial Attacks on Data Attribution

### v1

@@ Abstract
Data attribution aims to quantify the contribution of individual training data points to the outputs of an AI model, which has been used to measure the value of training data and compensate data providers. Given the impact on financial decisions and compensation mechanisms, a critical question arises concerning the adversarial robustness of data attribution methods. However, there has been little to no systematic research addressing this issue. In this work, we aim to bridge this gap by detailing a threat model with clear assumptions about the adversary’s goal and capabilities, and by proposing principled adversarial attack methods on data attribution. We present two such methods, Shadow Attack and Outlier Attack, both of which generate manipulated datasets to adversarially inflate the compensation. The Shadow Attack leverages knowledge about the data distribution in the AI applications, and derives adversarial perturbations through “shadow training”, a technique commonly used in membership inference attacks. In contrast, the Outlier Attack does not assume any knowledge about the data distribution and relies solely on black-box queries to the target model’s predictions. It exploits an inductive bias present in many data attribution methods—outlier data points are more likely to be influential—and employs adversarial examples to generate manipulated datasets. Empirically, in image classification and text generation tasks, the Shadow Attack can inflate the data-attribution-based compensation by at least 200%, while the Outlier Attack achieves compensation inflation ranging from 185% to as much as 643%.
@@ 1 Introduction
This work presents the first comprehensive study that aims to fill in this gap. We first outline the threat model by detailing the data compensation workflow and precisely specifying the assumptions we made about the adversary’s objective and capabilities. One key assumption is that the data contribution is periodic, and there is certain persistence across consecutive iterations of data attribution, which is the source of knowledge that the adversary could exploit. Motivating examples are provided to justify this assumption in practice. We also assume that the adversary may either have access to the distribution of the data used by the target AI model, or can get black-box queries of the target model’s predictions. Both assumptions are commonly seen in the AI security literature (Shokri et al., 2017; Chen et al., 2017).
Subsequently, we propose two adversarial attack strategies, Shadow Attack and Outlier Attack, respectively relying on different assumptions about the adversary’s capabilities. The Shadow Attack relies on the access to data distribution, and employs the “shadow training” technique commonly used in membership inference attacks (Shokri et al., 2017) to train “shadow models” that imitate the target model. The adversary can then directly perturb their dataset to achieve a higher compensation rate on these shadow models. The Outlier Attack, instead, does not assume knowledge about the data distribution but only relies on black-box queries of the target model’s predictions. The key idea behind this method lies in an inductive bias of many data attribution methods—outlier data points are more likely to be more influential. The proposed Outlier Attack utilizes adversarial examples (Goodfellow et al., 2015; Chen et al., 2017) to generate realistic outliers in a black-box fashion.
We conduct extensive experiments, including both image classification and text generation settings, to demonstrate the effectiveness of the proposed attack methods. Our results show that the Shadow Attack can inflate the adversary’s compensation to at least 200% and up to , while the Outlier Attack can inflate the adversary’s compensation to at least 185% and up to 456%. These results are achieved by adding imperceptible perturbations to real-world data features.
Overall, our study reveals a critical practical challenge in deploying data attribution methods for data valuation and compensation, i.e., the adversarial vulnerability of these methods. Moreover, the design of the proposed adversarial attack methods, especially the Outlier Attack that exploits a common inductive bias of data attribution methods, offers deeper insights into these vulnerabilities. These findings provide valuable directions for future research aimed at enhancing the robustness of data attribution methods.
@@ 2 Related Work
Membership inference attacks aim to infer whether a specific data point was used during the training of a machine learning model, typically without knowing the actual training dataset or having white-box access to the model (Shokri et al., 2017). The general strategy involves adversaries leveraging information such as model architecture, training data distribution, or black-box model predictions (Shokri et al., 2017; Song & Mittal, 2021). We refer the readers to Hu et al. (2022) for a detailed survey on this topic. One of the proposed attack methods, Shadow Attack, is inspired by the “shadow training” technique (Shokri et al., 2017) commonly used in membership inference attacks, where the adversary draws “shadow samples” following the same distribution as the actual training dataset used by the target model to be attacked, and trains “shadow models” based on the shadow samples to imitate the target model.
@@
Adversarial example (Goodfellow et al., 2015) is a well-known phenomenon where small, often imperceptible perturbations to input features can significantly alter the predictions of machine learning models, particularly deep neural networks. These perturbations can be generated through black-box queries to the model predictions (Chen et al., 2017; Ilyas et al., 2018; Guo et al., 2019). For a comprehensive review of adversarial examples, see the surveys by Yuan et al. (2019) and Chakraborty et al. (2021). In the proposed Outlier Attack, we employ black-box adversarial attack methods for generating adversarial examples to generate realistic outliers relative to the training dataset of the target model, without needing access to the training dataset or the model details.
@@ 3.1 The Data Compensation Scenario
We assume that the Data Providers contribute data to the AI Developer periodically, a common practice in many AI applications. For example, large language models need periodic updates to stay aligned with the latest factual knowledge about the world (Zhang et al., 2023); recommender systems must adapt to evolving user preferences (Zhang et al., 2020); quantitative trading firms rely on up-to-date information to power their predictive models [^1] See, for example, the Bloomberg market data feed: https://www.bloomberg.com/professional/products/data/enterprise-catalog/market/.; and generative models for music or art benefit from fresh, innovative works by artists to diversify their creative outputs (Smith et al., 2024).
However, such periodic data contribution introduces risks: a malicious Data Provider (referred to as an Adversary thereafter) could exploit information from previous iterations to adversarially manipulate their future data contribution, potentially inflating their compensation unfairly.
To formalize this scenario, without loss of generality, we consider two consecutive iterations of data contribution, denoted as time steps $t=0$ and $t=1$. At $t=0$, there is no Adversary and the training dataset consists solely of contributions from benign Data Providers. This dataset is represented as $Z_{0}\in\mathcal{Z}$, where $\mathcal{Z}$ is the set of all possible training datasets. At $t=1$, the training dataset is $Z_{1}=Z_{1}^{b}\cup Z_{1}^{a}$, where $Z_{1}^{b}\in\mathcal{Z}$ represents the training data provided by benign Data Providers, while $Z_{1}^{a}\in\mathcal{Z}$ is the set provided by the Adversary. We also assume that $Z_{1}^{b}\ \cap Z_{1}^{a}=\varnothing$, meaning there is no overlap between the two datasets.
For any dataset $Z\in\mathcal{Z}$, each data point $z\in Z$ is represented as a pair $z=(x,y)$, where $x\in\mathcal{X}$ is the input feature and $y\in\mathcal{Y}$ is the prediction target. Here, $\mathcal{X}$ and $\mathcal{Y}$ refer to the feature space and target space, respectively.
@@
Let $\mathcal{M}$ represent the AI model space, and let $\mathcal{T}\colon\mathcal{Z}\to\mathcal{M}$ denote a training algorithm, mapping a training dataset $Z\in\mathcal{Z}$ to a model $\mathcal{T}(Z)\in\mathcal{M}$.
In data attribution, we aim to understand how individual data points from a training dataset $Z\in\mathcal{Z}$ contribute to the model output on a target (validation) data point from a validation dataset $V$. Given $Z$ and $V$, a data attribution method derives a contribution function $\tau\colon Z\times V\to\mathbb{R}$ that assigns a real value $\tau(z,v)$ to each training data point $z\in Z$ for a given validation data point $v\in V$. Denote the set of all such function $\tau$’s as $\mathcal{C}$, and the set of all possible validation sets as $\mathcal{V}$. Therefore, a data attribution method can be formalized as a function $\mathcal{A}\colon\mathcal{Z}\times\mathcal{M}\times\mathcal{V}\to\mathcal{C}$.
@@ 3.2 The Adversary
We now discuss the objective and capabilities of the Adversary under the data compensation scenario outlined above.
@@
  • the exact training datasets $Z_{0}$ and $Z_{1}^{b}$;
  • the exact validation datasets $V_{0}$ and $V_{1}$;
@@
Assumption 2 (Access to Data Distribution and Training Algorithm)
@@
Assumption 3 (Black-box Access to Model)
Assume that the Adversary has access to black-box query to the model $\mathcal{T}(Z_{0})$, allowing the Adversary to query the model predictions on any input feature $x\in\mathcal{X}$.
Intuitively, 1 assumes persistence between the time steps $t=0$ and $t=1$, so that information gained from $t=0$ can inform the attack at $t=1$. In 2 and 3, it is worth noting that the Adversary’s access to information is limited to what is available at $t=0$ only.
@@
Finally, we define the action space available to the Adversary. We assume that the Adversary is restricted to making only small perturbations to an existing set of real data points in order to construct the adversarial dataset $Z_{1}^{a}$. This constraint implies that the Adversary cannot introduce entirely synthetic or arbitrary data, but rather, can modify real data points in a subtle manner, which reflects a realistic adversarial scenario where overly large or unnatural alterations would be easily detectable.
@@ 4.2 Adversarial Manipulation by Maximizing Shadow Compensation Rate
Given the shadow compensation rate, a natural idea is to apply adversarial perturbations to $Z$ by solving the following optimization problem
$Z_{1}^{a}=\operatorname*{arg\,max}_{Z^{\prime}\in\mathcal{N}(Z)}c_{s}(Z^{\prime}),$ (3)
where $\mathcal{N}(Z)$ specifies the space of datasets with undetectable perturbations.
In practice, however, solving this optimization problem has two technical challenges. Firstly, the objective $c_{s}(\cdot)$ is discrete and can be difficult to optimize. Secondly, when doing iterative-style optimization, such as gradient ascent, $\tau^{(i)}$’s need to be updated and re-evaluated at every step as it depends on $Z$ and hence $M^{(i)}=T(Z^{(i)}\cup Z)$. This leads to two problems: On the one hand, updating $M^{(i)}$’s requires retraining, which is computationally heavy; on the other hand, many data attribution methods are computationally expensive, even with all the data and (retrained) models available. Hence, solving the problem in Eq. 3 requires repeatedly retraining and running data attributions on perturbed datasets, which may be infeasible for many training algorithms and data attribution methods.
@@
which aims to directly maximize the contribution values of the data points in $Z$ on the shadow validation set $V^{(0)}$.
To address the second challenge, we first approximate the retrained models by the initial models trained with the original $Z$ (before any gradient ascent steps) for computational efficiency, and we further adopt Grad-Dot (Charpiat et al., 2019), one of the most efficient data attribution methods when evaluating the contribution function. [^3] Empirically, this works well even when the actual data attribution method used by the AI Developer is more advanced ones, as shown in Section 6.2. Using this method, the contribution value $\tau^{(i)}(z,v)$ for any training data point $z$ and validation data point $v$ equals the dot product between the gradients of the loss function on $z$ and that on $v$, evaluated with model $M^{(i)}$.
@@ 5 Outlier Attack
For large-scale AI applications, such as generative AI services, 2 might be overly strong, as it could be difficult for the Adversary to guess the distribution of the full training data. In this case, however, 3 often holds true.
In this section, we introduce another proposed attack method, Outlier Attack, which only relies on 1 and 3. See Figure 2 for an illustration.
@@ 5.2 Generating Realistic Outliers with Adversarial Attacks
Next, we propose a general strategy to address the two aforementioned challenges. To tackle the first challenge, we propose starting with a set of real-world data and transforming them into outliers through small adversarial perturbations. With careful design, this strategy also addresses the second challenge: if we can perturb a data point such that an AI model has low confidence in correctly predicting its target, then it is highly likely that this data point behaves as an outlier relative to the model’s training dataset. More specifically, we discuss two key design aspects for generating outliers by perturbing real data points.
@@
    Data component to be perturbed. Recall that a data point $z=(x,y)$ consists of both the input feature $x$ and the prediction target $y$. To obtain an outlier data point from $z$, we perturb the feature $x$ only, without modifying the target $y$. Intuitively, flipping the annotated target $y$ could easily result in an outlier data point through mislabeling, such an outlier typically degrades model performance and is likely to be identified as negatively influential. Moreover, mislabeled data points are easier for the AI Developer to detect. Therefore, it is more effective to perturb only the feature $x$ while keeping the target $y$ unchanged.
@@
Combining these two design choices, the resulting perturbed data points coincide with what are commonly referred to as adversarial examples in the literature of adversarial attacks against neural network models. Consequently, existing adversarial attack methods for deriving adversarial examples can be leveraged to generate realistic outliers in our problem setup. Notably, this strategy is very general and can be applied to a variety of data modalities and models by leveraging different off-the-shelf black-box adversarial attack methods.
@@
We consider two concrete machine learning settings, image classification, and text generation, and discuss the choices of adversarial attack methods. For smaller-scale image classification settings, we leverage Zeroth Order Optimization (ZOO) (Chen et al., 2017) based adversarial attacks, which approximates the gradient ascend on the loss function with respect to the data features using black-box queries to the target model $\mathcal{T}(Z_{0})$. For larger-scale image classification settings, we employ a more advanced black-box adversarial attack method, Simba (Guo et al., 2019), which is computationally more efficient. At a high level, Simba sequentially perturbs each scalar pixel value in the image by trying to perturb in both directions and accept the perturbation once it increases the loss. For the text generation setting, we utilize TextFooler (Jin et al., 2020), a black-box adversarial attack method tailored for text data. In all the methods of choice, the attack only requires the black-box queries to the predictions of the target model.

### v2

@@ Abstract
Data attribution aims to quantify the contribution of individual training data points to the outputs of an AI model, which has been used to measure the value of training data and compensate data providers. Given the impact on financial decisions and compensation mechanisms, a critical question arises concerning the adversarial robustness of data attribution methods. However, there has been little to no systematic research addressing this issue. In this work, we aim to bridge this gap by detailing a threat model with clear assumptions about the adversary’s goal and capabilities and proposing principled adversarial attack methods on data attribution. We present two methods, Shadow Attack and Outlier Attack, which generate manipulated datasets to inflate the compensation adversarially. The Shadow Attack leverages knowledge about the data distribution in the AI applications, and derives adversarial perturbations through “shadow training”, a technique commonly used in membership inference attacks. In contrast, the Outlier Attack does not assume any knowledge about the data distribution and relies solely on black-box queries to the target model’s predictions. It exploits an inductive bias present in many data attribution methods—outlier data points are more likely to be influential—and employs adversarial examples to generate manipulated datasets. Empirically, in image classification and text generation tasks, the Shadow Attack can inflate the data-attribution-based compensation by at least $200\%$, while the Outlier Attack achieves compensation inflation ranging from $185\%$ to as much as $643\%$.
@@ 1 Introduction
This work presents the first comprehensive study to fill this gap. We first outline the threat model by detailing the data compensation workflow and specifying the assumptions we made. One key assumption is that the data contribution is periodic, and there is certain persistence across consecutive iterations of data contributions, which is the source of knowledge that the adversary could exploit. We also assume that the adversary may either have access to the distribution of the data used by the target model of the AI system or can get black-box queries of the target model’s predictions, both are commonly seen in the AI security literature (Shokri et al., 2017; Chen et al., 2017).
Subsequently, we propose two adversarial attack strategies, Shadow Attack and Outlier Attack, respectively relying on different assumptions about the adversary’s capabilities. The Shadow Attack relies on the access to data distribution and employs the “shadow training” technique commonly used in membership inference attacks (Shokri et al., 2017) to train “shadow models” that imitate the target model. The adversary can then directly perturb their dataset to achieve a higher compensation on these shadow models. The Outlier Attack, instead, does not assume knowledge about the data distribution but only relies on black-box queries of the target model’s predictions. The key idea behind this method lies in an inductive bias of many data attribution methods—outlier data points are more likely to be more influential. The proposed Outlier Attack utilizes adversarial examples (Goodfellow et al., 2015; Chen et al., 2017) to generate realistic outliers in a black-box fashion.
We conduct extensive experiments, including both image classification and text generation settings, to demonstrate the effectiveness of the proposed attack methods. Our results show that by only adding imperceptible perturbations to real-world data features, the Shadow Attack can inflate the adversary’s compensation to at least $200\%$ and up to $456\%$, while the Outlier Attack can inflate the adversary’s compensation to at least $185\%$ and up to $643\%$.
Overall, our study reveals a critical practical challenge—adversarial vulnerability—in deploying data attribution methods for data valuation and compensation. Moreover, the design of the proposed attack methods, especially the Outlier Attack that exploits a common inductive bias of data attribution methods, offers deeper insights into these vulnerabilities. These findings provide valuable directions for future research to enhance the robustness of data attribution methods.
@@ 2 Related work
Membership inference attacks aim to infer whether a specific data point was used during the training of a machine learning model, typically without knowing the actual training dataset or having white-box access to the model (Shokri et al., 2017). The general strategy involves leveraging information such as model architecture, training data distribution, or black-box model predictions (Shokri et al., 2017; Song & Mittal, 2021). We refer the readers to Hu et al. (2022) for a detailed survey on this topic. One of the proposed attack methods, the Shadow Attack, is inspired by the “shadow training” technique (Shokri et al., 2017) commonly used in membership inference attacks, where the adversary draws “shadow samples” following the same distribution as the actual training dataset used by the target model to be attacked, and trains “shadow models” based on the shadow samples to imitate the target model.
@@
Adversarial example (Goodfellow et al., 2015) is a well-known phenomenon where small, often imperceptible perturbations to input features can significantly alter the predictions of machine learning models, particularly deep neural networks. These perturbations can be generated through black-box queries to the model predictions (Chen et al., 2017; Ilyas et al., 2018; Guo et al., 2019). For a comprehensive review of adversarial examples, see Yuan et al. (2019); Chakraborty et al. (2021). In the proposed Outlier Attack, we employ black-box adversarial attack methods for generating adversarial examples to generate realistic outliers relative to the training dataset of the target model, without needing access to the training dataset or the model details.
@@ 3.1 The data compensation scenario
We assume that the Data Providers contribute data to the AI Developer periodically, a common practice in many AI applications. For example, large language models need periodic updates to stay aligned with the latest factual knowledge about the world (Zhang et al., 2023); recommender systems must adapt to evolving user preferences (Zhang et al., 2020); quantitative trading firms rely on up-to-date information to power their predictive models [^1] See, for example, the Bloomberg market data feed: https://www.bloomberg.com/professional/products/data/enterprise-catalog/market/.; and generative models for music or art benefit from fresh, innovative works by artists to diversify their creative outputs (Smith et al., 2024). However, such periodic data contribution introduces risks: a malicious Data Provider (referred to as an Adversary thereafter) could exploit information from previous iterations to adversarially manipulate their future data contribution, potentially inflating their compensation unfairly.
To formalize this scenario, without loss of generality, we consider two consecutive iterations of data contribution, denoted as time steps $t=0$ and $t=1$. At $t=0$, there is no Adversary and the training dataset consists solely of contributions from benign Data Providers. This dataset is represented as $Z_{0}\in\mathcal{Z}$, where $\mathcal{Z}$ is the set of all possible training datasets. At $t=1$, the training dataset is $Z_{1}=Z_{1}^{b}\cup Z_{1}^{a}$, where $Z_{1}^{b}\in\mathcal{Z}$ represents the training data provided by benign Data Providers, while $Z_{1}^{a}\in\mathcal{Z}$ is the set provided by the Adversary. We also assume that $Z_{1}^{b}\ \cap Z_{1}^{a}=\varnothing$, meaning there is no overlap between the two datasets. Finally, for a dataset $Z\in\mathcal{Z}$, each data point $z\in Z$ is represented as a pair $z=(x,y)$, where $x\in\mathcal{X}$ is the input feature and $y\in\mathcal{Y}$ is the prediction target, with $\mathcal{X}$ and $\mathcal{Y}$ referring to the feature space and target space, respectively.
@@
Let $\mathcal{M}$ represent the set of AI models, and let $\mathcal{T}\colon\mathcal{Z}\to\mathcal{M}$ denote a training algorithm, mapping a training dataset $Z\in\mathcal{Z}$ to a model $\mathcal{T}(Z)\in\mathcal{M}$.
In data attribution, we aim to understand how individual data points from a training dataset $Z\in\mathcal{Z}$ contribute to the model output on a target (validation) data point from a validation dataset $V$. Given $Z$ and $V$, a data attribution method derives a contribution function $\tau\colon Z\times V\to\mathbb{R}$ that assigns a real value $\tau(z,v)$ to each training data point $z\in Z$ for a given validation data point $v\in V$. Denote the set of all such function $\tau$’s as $\mathcal{C}$, and the set of all possible validation sets as $\mathcal{V}$. Therefore, a data attribution method can be formalized as a function $\mathcal{A}\colon\mathcal{Z}\times\mathcal{M}\times\mathcal{V}\to\mathcal{C}$. In most cases, we will consider $\mathcal{A}(Z,\mathcal{T}(Z),V)$, hence $Z$ and $V$ alone suffice to specify the resulting $\tau$.
@@ 3.2 The adversary
We now discuss the Adversary’s objective and capability under the data compensation scenario.
@@
  • the exact training datasets ($Z_{0}$ and $Z_{1}^{b}$) and the exact validation datasets ($V_{0}$ and $V_{1}$);
@@
Assumption 2 (Access to data distribution and training algorithm).
@@
Assumption 3 (Black-box access to model).
Assume that the Adversary has access to black-box query to the model $\mathcal{T}(Z_{0})$ to query the model predictions on any input feature $x\in\mathcal{X}$.
Intuitively, 1 assumes persistence between time steps $t=0,1$, so that information gained from $t=0$ can inform the attack at $t=1$. In 2 and 3, it is worth noting that the Adversary’s access to information is limited to what is available at $t=0$ only.
@@
We assume that the Adversary is restricted to making only small perturbations to an existing set of real data points to construct the adversarial dataset $Z_{1}^{a}$. This implies that the Adversary cannot introduce entirely synthetic or arbitrary data, but rather, can modify real data points subtly. This reflects a realistic adversarial scenario as overly large or unnatural alterations would be easily detectable.
@@ 4.2 Adversarial manipulation by maximizing shadow compensation rate
Given the shadow compensation rate, a natural idea is to apply adversarial perturbations to $Z$ by solving $Z_{1}^{a}=\operatorname*{arg\,max}\nolimits_{Z^{\prime}\in\mathcal{N}(Z)}c_{s}(Z^{\prime})$, where $\mathcal{N}(Z)$ specifies the space of datasets with undetectable perturbations. In practice, however, solving this optimization problem has two technical challenges. Firstly, the objective $c_{s}(\cdot)$ is discrete and can be difficult to optimize. Secondly, when doing iterative-style optimization, such as gradient ascent, $\tau^{(i)}$’s need to be updated and re-evaluated at every step as it depends on $Z$ and hence $M^{(i)}=T(Z^{(i)}\cup Z)$. This leads to two problems: On the one hand, updating $M^{(i)}$’s requires retraining, which is computationally heavy; on the other hand, many data attribution methods are computationally expensive, even with all the data and (retrained) models available. Hence, maximizing $c_{s}(\cdot)$ requires repeatedly retraining and running data attributions on perturbed datasets, which may be infeasible for many cases.
@@
which aims to directly maximize the contribution values of the data points in $Z$ on the shadow validation set $V^{(0)}$. To address the second challenge, we first approximate the retrained models by the initial models trained with the original $Z$ (before any gradient ascent steps) for computational efficiency, and we further adopt Grad-Dot (Charpiat et al., 2019), one of the most efficient data attribution methods when evaluating the contribution function. [^3] Empirically, this works well even when the actual data attribution method used by the AI Developer is more advanced ones, as shown in Section 6.2. Using this method, the contribution value $\tau^{(i)}(z,v)$ for any training data point $z$ and validation data point $v$ equals the dot product between the gradients of the loss function on $z$ and that on $v$, evaluated with model $M^{(i)}$.
@@ 5 Outlier Attack
For large-scale AI applications such as generative AI services, 2 might be overly strong as it could be difficult for the Adversary to guess the distribution of the full training data. However, in this case, 3 often holds. In this section, we further propose Outlier Attack, which only relies on 1 and 3. See Figure 2 for an illustration.
@@ 5.2 Generating realistic outliers with adversarial attacks
To tackle the first challenge, we propose starting with a set of real-world data and transforming them into outliers through small adversarial perturbations. With careful design, this strategy also addresses the second challenge: if after the perturbation, an AI model has low confidence in correctly predicting its target, then this perturbed data point likely behaves as an outlier relative to the model’s training dataset. We discuss two key design aspects for achieving these goals:
@@
    Data component to be perturbed. Recall that a data point $z=(x,y)$ consists of both the input feature $x$ and the prediction target $y$. To obtain an outlier data point from $z$, we perturb the feature $x$ only, without modifying the target $y$. Intuitively, flipping the target $y$ could easily result in an outlier data point through mislabeling, such an outlier typically degrades model performance and is likely to be identified as negatively influential. Moreover, mislabeled data points are easier for the AI Developer to detect. Therefore, it is more effective to perturb only the feature $x$ while keeping the target $y$ unchanged.
@@
Combining these two design choices, the resulting perturbed data points coincide with what is commonly referred to as adversarial example in the literature of adversarial attacks against neural network models. Consequently, existing adversarial attack methods for deriving adversarial examples can be leveraged to generate realistic outliers in our problem setup. Notably, this strategy is very general and can be applied to a variety of data modalities and models by leveraging different off-the-shelf black-box adversarial attack methods.
@@
We consider two concrete machine learning settings, image classification, and text generation, and discuss the choices of adversarial attack methods. For smaller-scale image classification settings, we leverage Zeroth Order Optimization (ZOO) (Chen et al., 2017) based adversarial attacks, which approximates the gradient ascend on the loss function with respect to the data features using black-box queries to the target model $\mathcal{T}(Z_{0})$. For larger-scale image classification settings, we employ a more advanced black-box adversarial attack method, Simba (Guo et al., 2019), which is computationally more efficient. At a high level, Simba sequentially perturbs each scalar pixel value in the image by trying to perturb in both directions and accept the perturbation once it increases the loss. For the text generation setting, we utilize TextFooler (Jin et al., 2020), a black-box adversarial attack method tailored for text data. In all the methods of choice, the attack only requires black-box queries to get the predictions of the target model.
5.3 Theoretical understanding
The following theorem further provides theoretical insights about the effectiveness of the proposed Outlier Attack. The formal statement, notations, and the proof can be found in Appendix A.
Theorem 5.1 (Informal).
Consider a model trained by ERM on a dataset of size $n$ with a smooth loss $\ell$ with respect to model parameters $\theta$. Assume its corresponding influence score $\tau$, gradient $\nabla_{\theta}\ell(\theta,z)$, and Hessian $\nabla_{\theta}^{2}\ell(\theta,z)$ are all bounded, i.e., $|\tau|,\|\nabla_{\theta}\ell\|_{2},\|\nabla_{\theta}^{2}\ell\|_{\mathrm{op}}=\Theta_{n}(1)$. Assume the influence score $\tau$ is based on the influence function by Koh & Liang (2017), which takes the form
$\tau_{\text{IF}}(z_{j},z_{\text{test}};\hat{\theta})=-\nabla_{\theta}\ell(\hat{\theta},z_{\text{test}})^{\top}\left[\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}^{2}\ell(\hat{\theta},z_{i})\right]^{-1}\nabla_{\theta}\ell(\hat{\theta},z_{j}),$ (4)
where $z_{\text{test}}$ is a test data point while $\{z_{i}\}_{i=1}^{n}$ are the training data points and $\hat{\theta}$ is the model parameters trained on $\{z_{i}\}_{i=1}^{n}$. Then, for any $z_{\text{test}}$, when $z_{j}$ in $\{z_{i}\}_{i=1}^{n}$ is perturbed to $z_{j}^{\prime}$,
  • $\tau_{\text{IF}}(z_{i},z_{\text{test}};\hat{\theta}^{\prime})=\tau_{\text{IF}}(z_{i},z_{\text{test}};\hat{\theta})+O(1/n)$ for all $i\neq j$, and
  • $\tau_{\text{IF}}(z_{j}^{\prime},z_{\text{test}};\hat{\theta}^{\prime})=\tau_{\text{IF}}(z_{j}^{\prime},z_{\text{test}};\hat{\theta})+O(1/n)$,
where $\hat{\theta}^{\prime}$ refers to the model parameters trained on the perturbed dataset $\{z_{i}\}_{i\neq j}\cup\{z_{j}^{\prime}\}$.
Intuitively, perturbing $z_{j}$ through adversarial attack will increase the magnitude of the gradient $\nabla_{\theta}\ell(\hat{\theta},z_{j})$, which tends to also increase the influence score on the original model $\hat{\theta}$, i.e., $\tau_{\text{IF}}(z_{j}^{\prime},z_{\text{test}};\hat{\theta})\gg\tau_{\text{IF}}(z_{j},z_{\text{test}};\hat{\theta})$. However, in order to get a higher compensation share, the perturbed data point $z_{j}^{\prime}$ needs to have a high influence score on the model $\hat{\theta}^{\prime}$ trained on the perturbed dataset, i.e., $\tau_{\text{IF}}(z_{j}^{\prime},z_{\text{test}};\hat{\theta}^{\prime})$, which is not guaranteed by the adversarial attack without characterizing the new model $\hat{\theta}^{\prime}$. Theorem 5.1 exactly does this and asserts that $\tau_{\text{IF}}(z_{j}^{\prime},z_{\text{test}};\hat{\theta}^{\prime})$ will be close to $\tau_{\text{IF}}(z_{j}^{\prime},z_{\text{test}};\hat{\theta})$ when the datasets used to train $\hat{\theta}$ and $\hat{\theta}^{\prime}$ are close. It further guarantees that the influence scores of the rest unchanged data points will also remain similar. This explains why the effect of adversarial attacks on the original model can successfully translate to the new model. The results in Theorem 5.1 can be generalized to the case where more than one data point is perturbed, and a small set of clean data is also added to train the new model. The bound will change from $O(1/n)$ to $O(k/n)$, where $k$ is the total number of changed data points in the dataset.